# Nonlocal Locking of Observable Quantities: A Faithful Signature of Nonclassical Correlations

### v1

@@ Abstract
$\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})=\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\sigma_{\bf A})$, where $\sigma_{\bf A}=u_{1}\otimes u_{2}(\rho_{\bf A})u^{\dagger}_{1}\otimes u^{\dagger}_{2}$ for arbitrary $u_{i}$’s on the respective parts.
A simplified from of $\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ can be obtained for generic two-qubit state $\rho_{\bf A}=\frac{1}{4}(\mathbf{I}+\bm{\sigma}^{\intercal}_{1}{\bf r}_{1}+{\bf r}^{\intercal}_{2}\bm{\sigma}_{2}+\bm{\sigma}^{\intercal}_{1}{\bf T}\bm{\sigma}_{2})$, which is stated in the following theorem by considering $\mathbb{O}$ to be the Hamiltonian of the system.
@@
Proof is provided in Appendix VI. There we also obtain a closed letter expression of this quantity for the states having completely mixed marginals which is stated in the following Corollary.
@@
The plot depicted in Fig.1 exhibits the phenomena of energy locking in two-qubit Werner class of states. Notably, this phenomenon promises nontrivial usages of nonclassical correlation in devising efficient quantum batteries [55, 56, 57]. In multi-cell quantum batteries, stored energy can be lost due to various factors such as decoherence, global and local interactions, etc. For an isolated system the evolution is restricted to unitary processes only. Additionally, if interactions between individual systems are not allowed, energy drainage can occur only due to local interactions. Our findings indicate that energy storage is not possible in classically correlated states when local interactions and classical channels are available between pairs of isolated quantum batteries. In contrast, our Theorem 2 shows that for a system with a non-degenerate Hamiltonian, non-classical correlation in the form of discord can efficiently store energy in such a scenario.
Discussions.– We start by pointing out the recent work of Ref. [58]. There also the authors explored the role of nonclassical correlation in ergotropic work extraction. As established in their Theorem 1, the zero value of the ergotropic quantity implies the bipartite state to be classically correlated. But, the other way implication does not hold. Subsequently, this study is extended by considering generalised measurements [59], albeit the question of both way implication remains unresolved. In contrast, our Theorem 2 presents faithful quantifiers of nonclassical correlations for arbitrary bipartite systems. An interesting question is whether these quantities serve as valid measures of nonclassicality. Recall that, a correlation measure ($M$) is expected to fulfill the following three criteria [19]: (i) faithfulness, meaning $M(\rho_{\bf A})=0\Leftrightarrow\rho_{\bf A}$ is a CQ state, (ii) unitary invariance, indicating $M((u_{1}\otimes u_{2})(\rho_{\bf A})(u_{1}^{\dagger}\otimes u_{2}^{\dagger}))=M(\rho_{\bf A}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ u_{1}\in\mathrm{U}_{1}\leavevmode\nobreak\ \&\leavevmode\nobreak\ u_{2}\in\mathrm{U}_{2}$, and (iii) monotonicity, where $M((\mathcal{I}\otimes\Lambda)(\rho_{\bf A}))\leq M(\rho_{\bf A})$ holds for any CPTP map $\Lambda$ acting on the $A_{2}$ subsystem. While we have established that the first two properties are satisfied, the last one requires further study. Another intriguing question, which we leave for further exploration, is whether these nonclassical quantifiers are independent of the original quantum discord measure. Two quantifiers $M_{1}$ and $M_{2}$ are considered independent if there exist nonclassical states $\rho_{\bf A}$ and $\sigma_{\bf A}$ such that $M_{1}(\rho_{\bf A})>M_{1}(\sigma_{\bf A})$, while $M_{2}(\rho_{\bf A})<M_{2}(\sigma_{\bf A})$. Furthermore, although our current investigation primarily focuses on bipartite systems, the generality of our framework suggests potential for extension to multipartite systems.

### v3

@@ V Generic nonclassicality quantifiers and Ergodiscord
$\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})=\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\sigma_{\bf A})$, where $\sigma_{\bf A}=u_{1}\otimes u_{2}(\rho_{\bf A})u^{\dagger}_{1}\otimes u^{\dagger}_{2}$ for arbitrary $u_{i}$’s on the respective parts.
Proof.
While the Proposition holds for any observable $\mathbb{O}$, here we provide the proof by considering it to be the system’s Hamiltonian. Here we will use a property of ergodiscord which we prove for general observables in the Appendix (see equation (Proof.)).
$\mathbb{H}_{GN^{0}_{c}}(\rho_{{\bf A}})=\min_{U_{1}}E(\widetilde{\chi}_{\bf A}^{p})-E(\rho_{\bf A}^{p}),$ (1)
where $\rho^{p}_{\bf A}$ and $\widetilde{\chi}^{p}_{\bf A}$ respectively denotes the passive state corresponding to $\rho_{\bf A}$ and $N_{c}^{o}\circ U_{1}(\rho_{\bf A})$. Let $U_{1}^{\star}$ be the optimizing unitary operation. Then, the expression simplifies to
$\mathbb{H}_{GN^{0}_{c}}(\rho_{\bf A})=E(\chi_{\bf A}^{p})-E(\rho_{\bf A}^{p}),$ (2)
with $\chi^{p}_{\bf A}$ denoting the passive state corresponding to $N_{c}^{o}\circ U^{\star}_{1}(\rho_{\bf A})$. The evolution of the state is thus given by
[…]
Let $\sigma_{\bf A}=(V_{1}\otimes W_{2})\rho_{\bf A}(V_{1}^{\dagger}\otimes W_{2}^{\dagger})$, where $V_{1}$ and $W_{2}$ are some arbitrary local unitary on the $1$ and $2$ sub-parts respectively. The expression (2) in this case reads as,
$\mathbb{H}_{GN^{0}_{c}}(\sigma_{\bf A})=E(\eta_{\bf A}^{p})-E(\sigma_{\bf A}^{p}),$ (4)
and the state evolution [as describe in Eq.(3)] in this case becomes,
[…]
Note that $\sigma^{p}_{\bf A}=\rho^{p}_{\bf A}$ as $\rho_{\bf A}$ and $\sigma_{\bf A}$ are unitarily connected. Therefore, to establish the claim of the proposition, we only need to examine the invariance of $E(\eta^{p}_{\bf A})$ and $E(\chi^{p}_{\bf A})$. The proof proceeds by discussing these aspects in two parts.
  • P-I:
    Consider that $\sigma_{\bf A}=(I_{1}\otimes W_{2})\rho_{\bf A}(I_{1}\otimes W_{2}^{\dagger})$. The optimal state $\eta_{\bf A}$ achieved in Eq.(5) when starting with the initial state $\sigma_{\bf A}$ can also be achieved staring with the initial state $\rho_{\bf A}$. For that the second party applies the unitary $W_{2}$ on the state $\chi_{\bf A}$ in Eq.(3), i.e., $\eta_{\bf A}=(I_{1}\otimes W_{2})\chi_{\bf A}(I_{1}\otimes W_{2}^{\dagger})$, implying $E(\chi^{p}_{\bf A})=E(\eta^{p}_{\bf A})$.
  • P-I:
    Consider that $\sigma_{\bf A}=(V_{1}\otimes I_{2})\rho_{\bf A}(V_{1}^{\dagger}\otimes I_{2}^{\dagger})$. Assume that, $E(\chi^{p}_{\bf A})>E(\eta^{p}_{\bf A})$. On the other hand, we have
[…]
    This implies that starting from $\rho_{\bf A}$, they can reach to the passive state $\eta^{p}_{\bf A}$. This along with the assumption $E(\chi^{p}_{\bf A})>E(\eta^{p}_{\bf A})$ contradicts the claim that $U^{\star}_{1}$ in Eq.(3) is the optimal unitary, and thus $E(\chi^{p}_{\bf A})\ngtr E(\eta^{p}_{\bf A})$.
    Now, assume that $E(\chi^{p}_{\bf A})<E(\eta^{p}_{\bf A})$. Again this will contradict the optimality of $Z^{\star}_{1}$ in Eq.(5). Therefore we have, $E(\chi^{p}_{\bf A})=E(\eta^{p}_{\bf A})$.
(P-I) and (P-II) together thus implies $\mathbb{H}_{GN^{0}_{c}}(\rho_{\bf A})=\mathbb{H}_{GN^{0}_{c}}\big{(}(V_{1}\otimes W_{2})\rho_{\bf A}(V_{1}^{\dagger}\otimes W_{2}^{\dagger})\big{)}$. This completes the proof. ∎
VI Ergodiscord for general two-qubit states
A simplified from of $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ can be obtained for generic two-qubit state $\rho_{\bf A}=\frac{1}{4}(\mathbf{I}+\bm{\sigma}^{\intercal}_{1}{\bf r}_{1}+{\bf r}^{\intercal}_{2}\bm{\sigma}_{2}+\bm{\sigma}^{\intercal}_{1}{\bf T}\bm{\sigma}_{2})$.
@@ VI Ergodiscord for general two-qubit states
Proof.
A general 2-qubit state may be written as
$\rho=\frac{1}{4}\left[\mathbf{I}+\bm{\sigma}_{1}^{\intercal}\mathbf{r}_{1}+\mathbf{r}_{2}^{\intercal}\bm{\sigma}_{2}+\bm{\sigma}_{1}^{\intercal}\bm{T}\bm{\sigma}_{2}\right],$ (7)
where $\mathbf{r}_{1}$ is the Bloch vector of subsystem $A,$ $\bm{\sigma}_{1}=(\sigma_{x},\sigma_{y},\sigma_{z})_{A}$ is the vector of 3 Pauli matrices for subsystem $A$ (same idea for $B$) and $\bm{T}$ is the correlation matrix. These can be obtained via:
$\mathbf{r}_{1}={\sf Tr}\left[\rho\bm{\sigma}_{1}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{r}_{2}={\sf Tr}\left[\rho\bm{\sigma}_{2}\right],\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bm{T}={\sf Tr}\left[\rho\bm{\sigma}_{1}\bm{\sigma}_{2}^{\intercal}\right].$ (8)
Application of local unitary $U_{1}$ is equivalent to applying a 3x3 real rotation matrix to $\bm{\sigma}_{1}$, or equivalently replacing $\mathbf{r}_{1}\to\bm{R}\mathbf{r}_{1}$ and $\bm{T}\to\bm{R}\bm{T}$. Then we have
$\tau=U_{1}\rho U_{1}^{\dagger}=\frac{1}{4}\left[\mathbf{I}+\bm{\sigma}_{1}^{\intercal}\bm{R}\mathbf{r}_{1}+\mathbf{r}_{2}^{\intercal}\bm{\sigma}_{2}+\bm{\sigma}_{1}^{\intercal}\bm{R}\bm{T}\bm{\sigma}_{2}\right],$ (9)
The classical channel $\mathcal{N}^{0}_{C}$ between Alice and Bob is a measure and prepare channel. Any generic pair of qubit projectors can be written as $P_{\pm\bm{n}}=\ket{\bm{n}_{\pm}}\bra{\bm{n}_{\pm}}=\frac{1}{2}\left(\mathbf{I}\pm\bm{\sigma}_{1}^{\intercal}\bm{n}\right),$ where $\bm{n}$ is a unit vector in the Bloch sphere (it identifies the 2 orthogonal pure states forming the eigenbasis). This projective measurement also follows $\bra{\bm{n}_{\pm}}\bm{\sigma}_{1}^{\intercal}\ket{\bm{n}_{\pm}}=\pm\bm{n}^{\intercal}$. So the projected state of Bob after passing through $\mathcal{N}^{0}_{C}$ is
$\tau_{\pm}$ $=\bra{\bm{n}_{\pm}}\tau\ket{\bm{n}_{\pm}}_{A}$
$=\frac{1}{4}\left[\mathbf{I}\pm\bm{n}^{\intercal}\bm{R}\mathbf{r}_{1}+\left(\mathbf{r}_{2}^{\intercal}\pm\bm{n}^{\intercal}\bm{R}\bm{T}\right)\bm{\sigma}_{2}\right],$ (10)
The state accessible to Bob after all the above operations is $\chi=P_{\bm{n}}\tau_{\bm{n}}+P_{-\bm{n}}\tau_{-\bm{n}}$. We know that the eigenvalues of a qubit Hermitian operator of the form $B=b_{0}\mathbf{I}+\bm{b}^{\intercal}\bm{\sigma}_{2}$ are $\lambda_{\pm}=b_{0}\pm|\bm{b}|$. So we get the eigenvalues of the partial projections:
${\sf Spectrum}[\tau_{+\bm{n}}]$ =$=\left\{\frac{1}{4}\left[1+\bm{n}^{\intercal}\bm{R}\mathbf{r}_{1}\pm\left|\mathbf{r}_{2}^{\intercal}+\bm{n}^{\intercal}\bm{R}\bm{T}\right|\right]\right\},$
${\sf Spectrum}[\tau_{-\bm{n}}]$ =$=\left\{\frac{1}{4}\left[1-\bm{n}^{\intercal}\bm{R}\mathbf{r}_{1}\pm\left|\mathbf{r}_{2}^{\intercal}-\bm{n}^{\intercal}\bm{R}\bm{T}\right|\right]\right\}.$
Since the projectors $P_{\pm\bm{n}}$ are mutually orthogonal, the eigenvalues of $\chi$ follow directly from our previous results:
${\sf Spectrum}[\chi]=\left\{\!\begin{aligned} \frac{1}{4}\left[1+\bm{n}^{\intercal}\bm{R}\mathbf{r}_{1}+\left|\mathbf{r}_{2}^{\intercal}+\bm{n}^{\intercal}\bm{R}\bm{T}\right|\right],\\ \leavevmode\nobreak\ \frac{1}{4}\left[1+\bm{n}^{\intercal}\bm{R}\mathbf{r}_{1}-\left|\mathbf{r}_{2}^{\intercal}+\bm{n}^{\intercal}\bm{R}\bm{T}\right|\right]\\ \frac{1}{4}\left[1-\bm{n}^{\intercal}\bm{R}\mathbf{r}_{1}+\left|\mathbf{r}_{2}^{\intercal}-\bm{n}^{\intercal}\bm{R}\bm{T}\right|\right],\\ \leavevmode\nobreak\ \frac{1}{4}\left[1-\bm{n}^{\intercal}\bm{R}\mathbf{r}_{1}-\left|\mathbf{r}_{2}^{\intercal}-\bm{n}^{\intercal}\bm{R}\bm{T}\right|\right]\end{aligned}\right\}.$ (11)
We note that $\bm{R}$ always appears in the form $\bm{n}^{\intercal}\bm{R}$, so we may simplify the problem by defining a new unit vector ${\bf m}\equiv\bm{R}^{\intercal}\bm{n}$. Now, we only need to optimize over the vector ${\bf m}$, under the constraint $|{\bf m}|=1$, so this gives us
${\sf Spectrum}[\chi]=\{s_{ij}({\bf m}):=$
$\frac{1}{4}(1+(-1)^{i}{\bf m}^{\intercal}{\bf r}_{1}+(-1)^{j}|{\bf r}^{\intercal}_{2}+(-1)^{i}{\bf m}^{\intercal}{\bf T}|)\}_{i,j=0}^{1}.$ (12)
After arranging the spectrum in decreasing order we get $\{s_{ij}({\bf m})\}\mapsto\{q_{k}({\bf m})$}. Let $\{E_{k}\}$ be the non-degenerate increasing eigenvalues of $\mathbb{H}$, $\{p_{k}\}$ are the decreasing spectral of $\rho_{\bf A}$. Then the extractable ergotropy is.
$H_{GN^{0}_{c}}(\rho_{\bf A})$ $=\min_{u_{1}}E({\chi}_{\bf A}^{p})-E(\rho_{\bf A}^{p})$
$=\min_{m}\sum_{k=1}^{4}(q_{k}({\bf m})-p_{k})E_{k}.$ (13)
This completes the proof. ∎
We also obtain a closed letter expression of this quantity for the states having completely mixed marginals which is stated in the following Corollary.
@@
Proof.
A most general two-qubit state with maximally mixed marginals can be written as $\rho_{\bf A}=\frac{1}{4}\left[\mathbf{I}+\bm{\sigma}_{1}^{\intercal}\bm{T}\bm{\sigma}_{2}\right]$. This state can always be written as convex mixture of Bell state by applying local unitary $V_{1}\otimes W_{2}$, i.e.,
$\sigma_{\bf A}$ $=(V_{1}\otimes W_{2})\rho_{\bf A}(V_{1}^{\dagger}\otimes W_{2}^{\dagger})=\frac{1}{4}\left[\mathbf{I}+\bm{\sigma}_{1}^{\intercal}\bm{\tilde{T}}\bm{\sigma}_{2}\right]$
[…]
$\hskip 56.9055pt+p_{4}\ket{\psi^{-}}\bra{\psi^{-}},$
where $\bm{\tilde{T}}={\sf diag}(2(p_{1}+p_{3})-1,2(p_{2}+p_{3})-1),2(p_{1}+p_{2})-1)$. Denoting, $\omega=\max\{(p_{1}+p_{3}),(p_{2}+p_{3}),(p_{1}+p_{2})\}$, we have $\sqrt{\lambda_{\max}\left(\bm{T}\bm{T}^{\intercal}\right)}=2\omega-1$. Let us arrange the probabilities in decreasing order, $\{p_{i}\}\mapsto\{p^{\downarrow}_{i}\}$ with $p^{\downarrow}_{1}\geq p^{\downarrow}_{2}\geq p^{\downarrow}_{3}\geq p^{\downarrow}_{4}$. The energy of the $\sigma_{\bf A}^{p}$ and $\chi_{\bf A}^{p}$ are given by
$E(\sigma_{\bf A}^{p})=$ $\sum_{i=1}^{4}p_{i}^{\downarrow}E_{i},$
$E(\chi_{\bf A}^{p})=$ $\frac{1}{4}\left(1+\sqrt{\lambda_{\max}\left(\bm{T}\bm{T}^{\intercal}\right)}\right)(E_{1}+E_{2})$
$+\frac{1}{4}\left(1-\sqrt{\lambda_{\max}\left(\bm{T}\bm{T}^{\intercal}\right)}\right)(E_{3}+E_{4})$
$=$ $\frac{\omega}{2}(E_{1}+E_{2})+\frac{(1-\omega)}{2}(E_{3}+E_{4}).$
Thus we have,
$\mathbb{H}_{GN^{0}_{c}}(\rho_{\bf A})=\frac{\omega}{2}(E_{1}+E_{2})+\frac{1-\omega}{2}\left(E_{3}+E_{4}\right)-\sum_{i}p_{i}^{\downarrow}E_{i}.$
This completes the proof. ∎
Figure 3: Nonlocal energy locking in two qubit Werner states $\mathcal{W}^{\alpha}_{\bf A}$. Red dashed curve denotes the original quantum discord $\mathcal{Q}(\mathcal{W}^{\alpha}_{\bf A})=\frac{1-\alpha}{4}\log(1-\alpha)-\frac{1+\alpha}{2}\log(1+\alpha)+\frac{1+3\alpha}{4}\log(1+3\alpha)$ establishing the states are nonclassical for $\alpha\in(0,1]$ [17] (see also [58]). Blue curve amounts to the ergodiscord $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\mathcal{W}^{\alpha}_{\bf A})\leavevmode\nobreak\ [=\alpha\epsilon_{2}/2]$, i.e. the energy locked into the nonclassical correlation of the state $\mathcal{W}^{\alpha}_{\bf A}$. Here the plot is given by fixing $\epsilon_{2}=2$.
Using this result we derive and plot the ergodiscord for Werner class of state along with its quantum discord. Corollary 1, when applied to Werner class of states $\mathcal{W}^{\alpha}_{\bf A}=\alpha\phi^{max}_{\bf A}+(1-\alpha)\mathbf{I}_{\bf A}/4$, letting the Hamiltonian to be $h_{\bf A}=h_{A_{1}}\otimes\mathbf{I}_{A_{2}}+\mathbf{I}_{A_{1}}\otimes h_{A_{2}}$, with $h_{A_{1}}:=\epsilon_{1}\ket{1}\bra{1}\leavevmode\nobreak\ \&\leavevmode\nobreak\ h_{A_{2}}:=\epsilon_{2}\ket{1}\bra{1}$ |, we obtain $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\mathcal{W}^{\alpha}_{\bf A})=\alpha\epsilon_{2}/2$. Following facts are noteworthy:
The ergodiscord of the Werner state is independent of $\epsilon_{1}$, indicating that the energy difference between the two subsystems does not influence nonlocal energy locking. This is in stark contrast to the one-parameter family of states $\rho_{\bf A}^{\alpha}$ (see Figure 2), where increasing the energy gap between subsystems significantly alters the nature of energy locking.
From Figure 3, it is evident that for the Werner class of states, both quantum discord and ergodiscord increase monotonically with the parameter $\alpha$. However, this behavior does not hold for $\rho^{\alpha}_{\bf A}$. Figure 2 reveals that within certain ranges of $\alpha$, quantum discord exhibits monotonic growth, while ergodiscord decreases. Identifying the class of states for which the monotonicity of these two quantifiers aligns presents an intriguing avenue for further exploration.
VII Independence of ergodiscord and quantum discord
While both ergodiscord and quantum discord faithfully capture non-classicality of bipartite quantum states, our next result shows that they reflect independent notions.
Theorem 4.
There exist bipartite states $\rho_{\bf A},\sigma_{\bf A}\in\mathcal{D}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ such that, $\mathcal{Q}(\rho_{\bf A})<\mathcal{Q}(\sigma_{\bf A})$ but $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})>\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\sigma_{\bf A})$.
Proof.
Consider the one parameter family of states $\rho^{\alpha}_{\bf A}:=\alpha\phi^{\max}_{\bf A}+(1-\alpha)\ket{01}_{\bf A}\bra{01}$ |, with $\alpha\in[0,1]$, $\phi^{\max}_{\bf A}:=\ket{\phi^{\max}}_{\bf A}\bra{\phi^{\max}}$ |, and |$\ket{\phi^{max}}:=(\ket{00}+\ket{11})/\sqrt{2}$. Let system’s Hamiltonian be $h_{\bf A}=h_{A_{1}}\otimes\mathbf{I}_{A_{2}}+\mathbf{I}_{A_{1}}\otimes h_{A_{2}}$, with $h_{A_{1}}:=\epsilon_{1}\ket{1}\bra{1}\leavevmode\nobreak\ \&\leavevmode\nobreak\ h_{A_{2}}:=\epsilon_{2}\ket{1}\bra{1}$ |. As depicted in Fig. 2, for suitable choices of $\epsilon_{1},\epsilon_{2}$, the states possess the claimed feature. ∎
Not only does ergodiscord captures a distinct notion of nonclassicality as compared to the original quantum discord, but the set of free operations appropriate for bipartite quantum discord is also inadequate for it. Consider a bipartite system $\mathbf{A}$ composed of two subsystems $A_{1}$ and $A_{2}$. It is well established that, within the resource theory of quantum discord from $A_{1}\mapsto A_{2}$, the set of free operations must, at a minimum, include all completely positive trace-preserving (CPTP) maps acting on the $A_{2}$ subsystem [59]. However, this does not hold in the case of ergodiscord as established in the following.
Proposition 3.
For every $\alpha\in[0,1]$, there exist a bipartite quantum state |$\ket{\psi^{\alpha}}=\sqrt{1-\alpha/2}\ket{00}+\sqrt{\alpha/2}\ket{11}$ and a CPTP map $\Lambda^{\alpha}$ of the form
$\Lambda^{\alpha}\left(|0\rangle\langle 0|\right)=\frac{\alpha}{2-\alpha}|0\rangle\langle 0|+\frac{2(1-\alpha)}{2-\alpha}|1\rangle\langle 1|,$
$\Lambda^{\alpha}\left(|0\rangle\langle 1|\right)=\frac{\alpha}{\sqrt{\alpha(2-\alpha)}}|0\rangle\langle 1|,$
$\Lambda^{\alpha}\left(|1\rangle\langle 0|\right)=\frac{\alpha}{\sqrt{\alpha(2-\alpha)}}|1\rangle\langle 0|,$
$\Lambda^{\alpha}\left(|1\rangle\langle 1|\right)=|1\rangle\langle 1|,$
such that $\mathbf{I}_{A_{1}}\otimes\Lambda^{\alpha}_{A_{2}}(\psi^{\alpha}_{\bf A})=\rho^{\alpha}_{\bf A}$, where $\psi^{\alpha}_{\bf A}=\ket{\psi^{\alpha}}_{\bf A}\bra{\psi^{\alpha}}$ |.
Figure 4: The set of free operation of quantum discord and ergodiscord can not be the same. In the figure ergodiscord for two states $\psi^{\alpha}$ and $\rho_{\bf A}^{\alpha}$ has been plotted for Hamiltonian $h_{\bf A}=h_{A_{1}}\otimes\mathbf{I}_{A_{2}}+\mathbf{I}_{A_{1}}\otimes h_{A_{2}}$, with $h_{A_{1}}:=6\ket{1}\bra{1}\leavevmode\nobreak\ \&\leavevmode\nobreak\ h_{A_{2}}:=2\ket{1}\bra{1}$ |. The solid black curve represents $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho^{\alpha}_{\bf A})$ and the blue dotted curve corresponds to $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\psi^{\alpha}_{\bf A})$. Proposition 3 states that $\mathbf{I}_{A_{1}}\otimes\Lambda^{\alpha}_{A_{2}}(\psi^{\alpha}_{\bf A})=\rho^{\alpha}_{\bf A}$. From the figure it is easy to see that $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho^{\alpha}_{\bf A})\geq\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\psi^{\alpha}_{\bf A})\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \alpha$. Hence, local CPTP on $A_{2}$ part can increase the value of nonlocal energy locking from $A_{1}\mapsto A_{2}$. Therefore, unlike quantum discord, not all local CPTP on $A_{2}$’s part qualify as free operation of ergodiscord.
Proof.
Now, from Fig. 4, one can easily see that $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho^{\alpha}_{\bf A})\geq\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\psi^{\alpha}_{\bf A})\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \alpha$. It simply implies that the class of CPTP maps $\Lambda^{\alpha}$ can not be a free operation when one deals with ergodiscord from $A_{1}\mapsto A_{2}$. Whereas, in case of quantum discord from $A_{1}\mapsto A_{2}$ all CPTP maps on $A_{2}$ part, including all the $\Lambda^{\alpha}$’s, are free. This establishes the claim. ∎
Few observations as listed in the following remarks are in order:
Remark 1.
Local CPTP on $A_{2}$-part, however, can not increase energy locking from zero to non-zero. This follows from the fact that zero energy locking implies that the bipartite state is of the classical-quantum(CQ) form. If $\rho_{\bf A}$ be such CQ state, local CPTP on $A_{2}$-part can not make it a quantum-quantum or quantum-classical state. In other words, the set of all such CQ states is closed under any CPTP map on $A_{2}$ subsystem. This further highlights the faithfulness of ergodiscord.
Remark 2.
The ergodiscord of a bipartite system increases under local CPTP operations on $A_{2}$; however, unlike the geometric discord [60] based on Hilbert-Schmidt norm, this increase is inherently bounded. Specifically, the ergodiscord is always upper-bounded by the total energy of the system’s Hamiltonian. For instance, consider a two-qubit state $\rho_{\bf A}$ with Hamiltonians $h_{A_{1}}=\epsilon_{1}\ket{1}\bra{1}$ | and $h_{A_{2}}=\varepsilon_{2}\ket{1}\bra{1}$ |. After applying a local CPTP operation on subsystem $A_{2}$, the ergodiscord remains trivially bounded above by $\epsilon_{1}+\epsilon_{2}$. This avoids the unphysical behavior exhibited by geometric discord, whose increase is unbounded [59]. Moreover, while geometric discord is purely a mathematical construct with no clear physical interpretation, ergodiscord has a well-defined physical meaning that can be experimentally verified.
Remark 3.
The resource-theoretic formulations of ergodiscord and quantum discord are fundamentally different. Although both share the same set of free states, their free operations differ due to the distinct operational and physical contexts they address. A parallel can be drawn to the resource theories of coherence and asymmetry, which also share the same set of free states but differ in their free operations [61, 62, 63].
Remark 4.
One may argue that if the system Hamiltonian changes, the value of ergodiscord will also change for the same system. This can be understood as follows: although ergodiscord quantifies correlations, it implicitly depends on the observable (i.e., energy), and consequently, its value varies with the Hamiltonian. As a physically motivated quantity, ergodiscord is a byproduct of both the quantum state and the Hamiltonian. To address this arbitrariness, one might normalize ergodiscord by dividing it by the maximum energy eigenvalue minus the minimum energy eigenvalue of the Hamiltonian ($[\epsilon^{A}_{\max}+\epsilon^{B}_{\max}]-[\epsilon^{A}_{\min}+\epsilon^{B}_{\min}]$). This normalization ensures that the quantity remains zero for non-discordant states and finite for arbitrary Hamiltonians in the presence of discord. From a physical perspective, this normalized measure would represent the ratio of energy locked in correlations to the maximum ergotropy.
VIII Conclusion
In this work, we build an elegant operational approach to study nonclassical correlations. Our approach relies on the different probing method one can apply on a system consisting of specially separated subsystems. The success of our approach lies in the realization that when specially separated systems are correlated, that very correlation can lock away the values of observable quantities from an observer who is constrained to restricted forms of probing mechanisms. The power of this simple fact is exhibited in Observation 1. However, the scenario becomes a bit tricky once one aims to study nonclassical correlations. One needs to properly define what ‘classical probing’ means in the context of our approach, which mathematically is nothing but a subset of classical channels shared between the spatially separated sub-parts. Correctly identifying the subset relies on the observable one wants to measure (please refer to Fig. 5 in the appendix). If a resource theoretic formalism of the same is available, the job is relatively simpler. However, we go on to show the generality of our approach by building on the framework for observables (ergotropy) which, to the best of our knowledge, does not have any resource theory of its own. We term this nonclassicality as ergodiscord and observed an intriguing phenomenon called nonlocal locking of energy.
The phenomenon of energy locking, apart from its foundational importance, highlights the nontrivial utility of nonclassical correlations in designing efficient quantum batteries [21, 22, 23]. In multi-cell quantum batteries, stored energy can be lost due to factors such as decoherence and interactions, both local and global. For isolated systems evolving under unitary processes, energy drainage occurs solely through local interactions when inter-system interactions are absent. Our findings demonstrate that classically correlated states cannot store energy when only local interactions and classical communication channels are available between pairs of isolated quantum batteries. Theorem 2 establishes that for systems with non-degenerate Hamiltonians, nonclassical correlations quantified by ergodiscord enable efficient energy storage under these constraints. Moreover, Theorem 4 reveals the intriguing result that mixed states can store more energy than the maximally entangled state of the same system. Importantly, the notion of ergodiscord is different than the concept proposed in [35], which in literature is also known as thermal discord [64]. Recent works, such as Refs. [65, 66], have also examined the role of nonclassical correlations in ergotropic work extraction. However, the measures used in those studies are not faithful indicators of nonclassical correlations. In contrast, ergodiscord provides a reliable measure with clear operational significance. While our analysis primarily focuses on bipartite systems, the generality of our framework opens avenues for exploring energy storage in multipartite quantum systems.